OpenStudy (anonymous):

Given the triangle and equation below, what is the value of x? csc∠A + tan∠C = 5x

OpenStudy (anonymous):

wheres the triangle?

OpenStudy (anonymous):

do you know the answer ?

OpenStudy (anonymous):

right you can find the length of side AB by pythagoras ab = sqrt(7^2 - 6^2) = sqrt13 csc A = AC/AB = 7 / sqrt13 tan C = AB / BC = sqrt13 / 6 so to find x you solve : 7/sqrt13 + sqrt13 / 6 = 5x

OpenStudy (anonymous):

can you complete this?

OpenStudy (anonymous):

so is it 7+sqrt13 / 30?

OpenStudy (anonymous):

lets see we can multiply through by the LCM 6 sqrt13: 42 + 13 = 30sqrt13x x = 55 / 30sqrt13 = 11 / 6sqrt13

OpenStudy (anonymous):

- which is 0.508 to 3 dec places

OpenStudy (anonymous):

do you follow my solution mis.kayla?

OpenStudy (anonymous):

we can solve a fractional equation by multiplying every term by the LCM: 3 2 - + -- = 8 5x 3 LCM = 15x 3 * 15x 2 * 15x ------ + ------ = 15x*8 5x 3 simplifying: 3*3 + 2 * 5x = 120x